# PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling

### v1

@@ 4.4 Experimental Details
QG
Text Quality Diversity
GPT-2 1 124.8 0.141 0.759 4.564 3.130 0.176 0.210 0.629
DiffuSeq 20 395.0 0.149 0.730 1.555 1.274 0.901 0.170 0.564
DINOISER 2000 155.9 0.159 0.776 1.396 1.121 0.944 0.166 0.553
DiffusionBert 2000 513.6 0.150 0.712 3.040 2.209 0.566 0.392 0.759
Diffusion-EAGS 5 80.7 0.121 0.782 4.646 3.538 0.152 0.403 0.798
QQP
GPT-2 1 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
DiffuSeq 2000 124.247 0.00674 0.709 1.927 1.242 0.813 0.226 0.543
DINOISER 20 79.742 0.0042 0.821 1.421 1.126 0.935 0.264 0.542
Diffusion-EAGS 5 48.106 0.683 0.824 4.006 2.390 0.338 0.421 0.832
Table 2: QG & QQP Generation
@@
Table 3: Social Generation
@@ 5 Results
As shown in Table 234, our model consistently exhibits exceptional performance in terms of text quality while simultaneously maintaining diversity when compared to baseline models. As illustrated in Table 2, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP. In Table 3 Paradetox, our model demonstrates superior performance across all evaluated metrics. Such phenomenon represents that our model based on encoder-based PLMs show robustness on diverse perturbations from daily dialogues. When PPL exceeds 1,000, the model is considered to have failed in generating natural sequences, and is thus represented in gray color. In the context of Deontology, our model exceeds the baseline PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score of 4.755. Additionally, in Table 4 where semantic consistency is crucial, our model significantly outperforms the original dataset’s PPL by 23 points by leveraging PLMs while maintaining a high diversity score of 4.837.
5.1 Ours vs AR model
Diffusion-EAGS demonstrates high level of text quality surpassing that of GPT-2 in Table 2 in text quality. Notably, our model effectively reflects the pattern and style of the dataset, attributed to the capability of encoder-based PLMs. Specifically, as reported in Table 3-ParaDetox, our model excels at capturing semantic information with the highest MAUVE score of 0.733, despite ParaDetox being a challenging colloquial dataset including slang, numerous abbreviations, and various perturbations. In addition, our model demonstrate robustness to such perturbations with 69.5 PPL, while GPT-2 shows a lower performance of 389.1 PPL.
As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23, and 4. These results underscore our model’s efficacy in generating diverse responses.
@@ 5.1 Ours vs AR model
Table 4: ROC Generation
Specifically, in the context of story generation as shown in Table 4, our model not only demonstrates effectiveness in enhancing both the quality and diversity of the text but also maintains coherence with the contextual storyline. This is further supported by the high LLM-c scores, which confirm the plausibility of the narratives produced by our model.
5.2 Ours vs CDLMs
Notably, CDLMs demonstrate a noticeable deficiency in diversity. In contrast, our model excels at producing significantly more diverse sequences. This enhanced performance can be attributed to the integration of PLM into our framework, which evidently enriches the diversity of the generated outputs. In addition, our model generally achieves higher text quality scores than CDLMs across all experiments. Diffusion-EAGS not only surpasses Diffuseq and DINOISER in PPL and SOME but also achieves a MAUVE score improvement of over 0.4, thus indicating that the generated responses accurately reflect the dataset while preserving grammatical integrity. Furthermore, our models require only a few steps, resulting in a 40% reduction of costs with higher quality and diversity.
5.3 Ours vs DDLMs
Contrary to CDLMs, DiffusionBERT exhibits limitations in text quality, as shown in Table 234. On the other hand, our model achieves significantly higher scores across all quality metrics. This observation suggests that the naive application of BERT within the diffusion process fails to fully harness the capabilities of PLMs. SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox. Even in open-ended generation settings of Table 4, Diffusion-EAGS shows higher text quality and semantic scores. Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach. Above results further exemplify the high performance and the time efficiency of our model.
@@ 6.1 Ablation Study
The omission of EAGS initially results in a significant decline in performance, a noticeable decrease in text quality and diversity.
@@
Table 5: Ablation Study
We find that EAGS significantly contributes to a gradual entropy decrease, as shown in Appendix H. These observations highlight the critical role of selecting the sampling position based on given information in sequence generation. Additionally, excluding the use of the diffusion generation process, without the stepwise sampling, leads to a drastic reduction in overall performance, with PPL increasing by more than 1000 points and diversity scores dropping below 2. Consequently, our cMRF approach, integrated with the diffusion model, proves indispensable in aiding the PLM in effectively generating sequences. Furthermore, it is evident that without the incorporation of PLMs, the generation of natural sequences is unachievable.
@@ 6.2 Case Study: Keyword Based Generation
Our model operating within discrete space enables us to manipulate the output sequences using explicit instructions. To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences. Initially, we provide the same contextual input while varying the keywords. In the masked states, we randomly select positions, replacing them with the specified keywords. The results in Table 7 demonstrate that the generated sequences seamlessly integrate the keywords with context-specific semantics.
@@ 6.3 Bilinguality & Low Resource
Table 6: Translation Results
@@
We establish baselines using non-autoregressive (NAR) transformer models, such as Mask-and-Predict (Ghazvininejad et al., 2019a) and Easy-First (Kasai et al., 2020) - which bear resemblance to discrete diffusion models - diffusion models such as Diffuseq-v2 and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB. For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023).
Table 6 shows that predicting the target sequence without leveraging a multilingual PLM proves to be challenging. All diffusion baseline models struggle to produce correct outputs. For example, the pretrained SEDD model fails to effectively leverage conditional information, even after finetuning on the training datasets, consistent with the limitation observed in Section 5.3. Similar challenges arise in NAR transformer baselines. Despite constructing the vocabulary using the pretrained mBART-50 model (DisCo-m), the underlying issues remain. On the other hand, our proposed model, by incorporating a PLM, demonstrates promising results.
Interestingly, the output of the pretrained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token. While such issues result in lower BLEU scores, COMET and XCOMET often interpret them as semantically coherent, indicating a potential direction for future work to improve translation evaluation metrics. Despite these phenomena, a performance gap between translation models and DDLM remains. This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap. Details are in Appendix G.
@@ 6.4 Diversity Saturation
Inspired by the observation that Diffusion-EAGS consistently excel in terms of diversity across all results, we delve further into the diversity capabilities of our model. We assess the diversity performance in dataset-guided generation compared to LLMs. We measure the VS for 5 to 100 generations under a single condition. Such experiment demonstrates the extent to which the model’s output diversity saturates, enabling a comparison of asymptotic diversity performance. The experiment is conducted on the ‘deontology’ dataset which allows high output diversity in its settings. Details of using LLMs are provided in Appendix E.
@@ 7 Related Works
Continuous diffusion models Diffusion-LM (Li et al., 2022) propose a method of mapping tokenized sequences to embedding dimensions guided by a pretrained classifier. DiffuSeq-v1, v2 (Gong et al., 2023a, b) apply partial noising techniques. The core of these diffusion methodologies lies in the addition and restoration of random noise to facilitate generation. However, Cold Diffusion (Bansal et al., 2022) argues that such operations do not necessarily have to be governed by stochastic randomness.
Discrete diffusion models Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al. (2022) adopt PLMs to DDLM. SEDDLou et al. (2024) proposes score entropy inspired by MLM loss. Recent works by Shi et al. (2024) and Sahoo et al. (2024) extend this idea to propose a simplified view of the discrete framework, and obtain better empirical results. Zheng et al. (2024) further closes the gap between discrete diffusion models and masked models. These researches make an improvement on the open ended generation task. Venkatraman et al. (2024) showcase that SEDD priors can be used to sample from a posterior distribution, namely text infilling.
@@ 8 Conclusions
In this work, we introduce Diffusion-EAGS, integrating encoder-based PLMs with diffusion models for dataset-guided generation. By leveraging step-wise diffusion models and the property of MLM, we demonstrate that MLM-based PLMs can serve as denoising function with Entropy-based Adaptive Gibbs Sampling. We also propose entropy-based noise scheduling during training for adaptive sampling. Experimental results show that Diffusion-EAGS outperforms existing baselines, achieving high diversity and precise token-level control. These results demonstrate its enhanced text generation quality and diversity, as well as wide applicability.
@@
While Diffusion-EAGS demonstrates significant improvements in conditional generation tasks, there are several limitations. First, the quality of the text generated by our model depends on PLMs such as BERT. Second, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored. Future research could explore extending this method to other NLP tasks, as well as investigating the potential for reducing computational costs through more efficient entropy calculation techniques. Additionally, addressing the biases inherent in pretrained models remains a critical area of focus to ensure the robustness and fairness of generated outputs.
@@
Appendix A BERT as a MRF
@@
$\log\phi_{l}(X)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}),&\text{if }\text{[M]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$ (3)
@@
Appendix B Dataset Explanations
@@
Appendix C Experimental Outputs
C.1 LLM Evaluation
@@ C.2 Tables
Appendix D Well-Generated Output Examples
D.1 QG
Table 15: QG generation examples
QG Generation Output Examples of Diffusion-EAGS
Output $X_{1}$ “What kind of bird can fly backwards?”
Output $X_{2}$ “Which bird is able to fly backwards?”
Output $X_{1}$ “What is a pouch?”
Output $X_{2}$ “What is the smallest animal without a pouch.”
D.2 QQP
Table 16: QQP generation examples
QQP Generation Output Examples of Diffusion-EAGS
Output $X_{2}$ “What are the best known fiction and books of Isaac Asimov?”
Constraint $Y$ “Can we ever store energy produced in lightning?”
Output $X_{1}$ “How do we store heat energy from lightning?”
Output $X_{2}$ “How can you store energy from lightning?”
D.3 ParaDetox

### v2

@@ 4.4 Experimental Details
CMLM w/ Mask-Predict 10 669.9 0.0234 0.588 1.000 1.000 1.000 0.451 0.633
DisCo w/ Easy-First 10 716.1 0.0344 0.576 1.000 1.000 1.000 0.438 0.583
DiffusionBert 2000 775.9 0.737 0.716 3.101 2.058 0.599 0.424 0.826
@@
Table 2: Social Generation – Diversity values associated with higher perplexity (PPL) are displayed in gray, as increased perplexity typically indicates degenerate sequences.
QQP
GPT-2 1 66.270 0.112 0.754 3.886 2.566 0.423 0.344 0.787
DiffuSeq 2000 124.247 0.00674 0.709 1.927 1.242 0.813 0.226 0.543
DINOISER 20 79.742 0.0042 0.821 1.421 1.126 0.935 0.264 0.542
Diffusion-EAGS 5 48.106 0.683 0.824 4.006 2.390 0.338 0.421 0.832
QG
Text Quality Diversity
GPT-2 1 124.8 0.141 0.759 4.564 3.130 0.176 0.210 0.629
DiffuSeq 20 395.0 0.149 0.730 1.555 1.274 0.901 0.170 0.564
DINOISER 2000 155.9 0.159 0.776 1.396 1.121 0.944 0.166 0.553
DiffusionBert 2000 513.6 0.150 0.712 3.040 2.209 0.566 0.392 0.759
Diffusion-EAGS 5 80.7 0.121 0.782 4.646 3.538 0.152 0.403 0.798
Table 3: QG & QQP Generation
@@ 5 Results
As shown in Table 234, our model consistently exhibits exceptional performance in terms of text quality while simultaneously maintaining diversity when compared to baseline models.
In Table 2 Paradetox, our model demonstrates superior performance across all evaluated metrics. Such phenomenon represents that our model based on encoder-based PLMs show robustness on diverse perturbations from daily dialogues. When PPL exceeds 600, the model is considered to have failed in generating natural sequences, and is thus represented in gray color. In the context of Deontology, our model exceeds the baseline PPL and MAUVE scores, whereas SOME score represent the sufficient quality of text with the highest diversity score of 4.755. In the paradetox experiment, the text quality produced by the CMLM and SEDD models was found to be low. Consequently, these models were excluded from subsequent experiments. As illustrated in Table 3, Diffusion-EAGS generates the responses with the highest PPL score for QG, and highest MAUVE and PPL score for QQP. Additionally, in Table 4 where semantic consistency is crucial, our model significantly outperforms the original dataset’s PPL by 23 points by leveraging PLMs while maintaining a high diversity score of 4.837.
Diffusion-EAGS demonstrates the highest MAUVE score of 0.733 in Table 2-ParaDetox, and high level of text quality surpassing that of GPT-2 in Table 3 in text quality. ParaDetox is colloquial dataset including slang, numerous abbreviations, and various perturbations, so our model demonstrate robustness to such perturbations with 69.5 PPL. As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23, and 4. These results underscore our model’s efficacy in generating diverse responses.
@@
Table 4: ROC Generation – Because ROC is open-ended generation task, we include SEDD in this experiment.
Notably, CDLMs demonstrate a noticeable deficiency in diversity. In contrast, our model excels at producing significantly more diverse sequences. Furthermore, our models require only a few steps, while resulting in higher quality and diversity.
DiffusionBERT exhibits limitations in text quality, as shown in Table 234. On the other hand, our model achieves significantly higher scores across all quality metrics. This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs. SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox. Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach. A detailed comparison is provided in Appendix D.
@@ 6.1 Ablation Study
The omission of EAGS initially leads to a substantial decline in performance and text quality, producing degenerated results similar to those of traditional CMLMs.
@@
Table 5: Ablation Study : Deont denotes Deontology
We find that EAGS significantly contributes to a gradual entropy decrease, as shown in Appendix J. These observations highlight the critical role of selecting the sampling position based on given information in sequence generation. Additionally, excluding the use of the diffusion generation process, without the stepwise sampling, leads to a drastic reduction in overall performance, with PPL increasing by more than 1000 points and diversity scores dropping below 2. Consequently, our cMRF approach, integrated with the diffusion model, proves indispensable in aiding the PLM in effectively generating sequences. Furthermore, it is evident that without the incorporation of PLMs, the generation of natural sequences is unachievable.
@@ 6.2 Case Study: Keyword Based Generation
Our model operating within discrete space enables us to manipulate the output sequences using explicit instructions. To further explore this capability, we conduct the generation of sequences based on keywords positioned in the middle and at the end of masked sequences, which is challenging for AR models (Keskar et al., 2019). Initially, we provide the same contextual input while varying the keywords. In the masked states, we randomly select positions, replacing them with the specified keywords. The results in Table 7 demonstrate that the generated sequences seamlessly integrate the keywords with context-specific semantics.
@@ 6.3 Bilinguality & Low Resource Settings
Table 6: En-De Translation Results
@@
Although our baseline diffusion models demonstrated significant performance improvements over CMLMs, CMLMs were primarily developed to enhance translation performance. Therefore, we conducted additional experiments to further validate our methodology such as Mask-and-Predict and Easy-First, diffusion models such as Diffuseq-v2 (Gong et al., 2023b) and SEDD, traditional translation models such as mBART-50 (Tang et al., 2020) and NLLB. For evaluation metrics, we utilize sacreBLEU (Post, 2018) and neural-net scores such as COMET (Rei et al., 2020) and XCOMET (Guerreiro et al., 2023).
Table 6 shows that predicting the target sequence without leveraging a multilingual PLM proves to be challenging. All diffusion baseline models struggle to produce correct outputs. For example, the pretrained SEDD model fails to effectively leverage conditional information, even after finetuning on the training datasets, consistent with the limitation observed in Section D.3. Similar challenges arise in NAR transformer baselines. Despite constructing the vocabulary using the pretrained mBART-50 model (DisCo-m), the underlying issues remain. On the other hand, our proposed model, by incorporating a PLM, demonstrates promising results. In addition, these results show that our methodology can also be used with encoder-decoder models.
Interestingly, the output of the pretrained NLLB model (NLLB-naive-600M, not finetuned) reveal that neural network-based metrics are susceptible to the interference problem, specifically translated by other languages, even though we provide the language specific token. While such issues result in lower BLEU scores, COMET and XCOMET often interpret them as semantically coherent, indicating a potential direction for future work to improve translation evaluation metrics. Despite these phenomena, a performance gap between translation models and DDLM remains. This suggests that future research should address the semantic capabilities of diffusion models to help bridge this gap. Details are in Appendix I.
@@ 6.4 Diversity Saturation
Inspired by the observation that Diffusion-EAGS consistently excel in terms of diversity across all results, we delve further into the diversity capabilities of our model. We assess the diversity performance in conditional generation compared to LLMs. We measure the VS for 5 to 100 generations under a single condition. Such experiment demonstrates the extent to which the model’s output diversity saturates, enabling a comparison of asymptotic diversity performance. The experiment is conducted on the ‘deontology’ dataset which allows high output diversity in its settings. Details of using LLMs are provided in Appendix G.
@@ 7 Related Works
Continuous diffusion models Diffusion-LM (Li et al., 2022) propose a method of mapping tokenized sequences to embedding dimensions guided by a pretrained classifier. DiffuSeq-v1, v2 (Gong et al., 2023a, b) apply partial noising techniques. The core of these diffusion methodologies lies in the addition and restoration of random noise to facilitate generation. However, authors of Cold Diffusion (Bansal et al., 2022) argue that such operations do not necessarily have to be governed by stochastic randomness.
Discrete diffusion models Building on this rationale, D3PM (Austin et al., 2023) propose the discrete restoration-generation approach and DiffusionBERT He et al. (2022) adopt PLMs to DDLM. SEDDLou et al. (2024) propose score entropy inspired by MLM loss. Recent works by Shi et al. (2024) and Sahoo et al. (2024) extend this idea to propose a simplified view of the discrete framework, and obtain better empirical results. Zheng et al. (2024) further close the gap between discrete diffusion models and masked models, and also, by correcting the numerical precision error during the sampling process in SEDD-based models, reveal that the true generation perplexity of DDLMs lags behind that of AR models. These research make an improvement on the open ended generation task. Furthermore, Venkatraman et al. (2024) showcase that SEDD priors can be used to sample from a posterior distribution, namely text infilling, and Nie et al. (2024) demonstrate that DDLMs are scalable, and have advantages over traditional AR models.
@@ 8 Conclusions
In this work, we introduce Diffusion-EAGS, integrating PLMs with diffusion models for conditional generation. By leveraging step-wise diffusion models and the property of MLM, we demonstrate that MLM-based PLMs can serve as denoising function with Entropy-based Adaptive Gibbs Sampling. We also propose entropy-based noise scheduling during training for adaptive sampling. Experimental results demonstrate that Diffusion-EAGS outperforms existing baselines, yielding improved text generation quality, enhanced diversity, broad applicability, and more precise token-level control.
@@
While Diffusion-EAGS demonstrates significant improvements in conditional generation tasks, there are several limitations. Firstly, as our method is currently focused on text generation tasks, its applicability to text classification tasks, such as Named Entity Recognition and Part-of-Speech Tagging, remains unexplored. Future research could explore extending this method to other NLP tasks. Secondly, future works could investigate the potential in reducing computational costs through more efficient entropy calculation techniques. Lastly, although our current efforts concentrate on developing and validating our framework using encoder-only and encoder-decoder architectures, the potential integration of autoregressive (AR) models remains unexplored. However, since the denoising function—with its entropy-based, sequential ordering—aligns well with autoregressive decoding strategies, we posit that AR models can also serve as effective initializations within our proposed framework.
@@
Appendix A Measuring Potential Function $\phi(X)$ in MLM
In this section, we provide additional experimental details and results to support the observation that open-ended Masked Language Models (MLMs) exhibit increased potentials for the same sequence under different dataset constraints.
A.1 Experimental Setup
  • Model We use the pre-trained BERT large model (bert-large-cased) as the base model for all experiments. Additionally, we incorporate RocStories-conditioned guidance with the pre-trained model and use a fine-tuned BERT model on the RocStories dataset to evaluate the impact of dataset-specific constraints.
  • Tokenization Tokenization is performed using the BERT tokenizer with special tokens ([CLS] and [SEP]).
  • Potential Calculation The potential function $\phi(X)$ is computed as described in Equation 1 and  3 , where the log-potentials are obtained for each token using masked token logits.
  • Datasets
    • RocStories: Structured narratives from the RocStories dataset.
A.2 Results of Experiment and Implications for Conditional Generation
Using the BERT-large-cased model, the average log potential value for the standard MLM was 156.6150, while incorporating RocStories guidance increased this value to 175.5332, highlighting the impact of dataset-specific constraints. Additionally, fine-tuning the same model on RocStories resulted in an average potential function value of 3.7551 (on an exponential scale), demonstrating substantial variation introduced by conditional generation settings.
The results demonstrate the significant influence of dataset structure on the potential function in MLMs. Specifically, structured datasets like RocStories enforce stronger narrative constraints, leading to higher potentials and greater coherence in sequence generation. This supports the main text’s argument that conditionally constrained spaces enhance the consistency and predictability of MLM outputs.
Appendix B BERT as a MRF
@@
$\log\phi_{l}(X)=$$\begin{cases}1h(x_{l})^{T}f_{\theta}(X\backslash\{x_{l}\}),&\text{if }\text{[MASK]}\notin X_{l-1:l+1}\\ \\ 0,&\text{otherwise}\end{cases}$ (4)
@@
Appendix C Dataset Explanations
@@
Appendix D Detailed analysis of Results
D.1 Ours vs AR model
Notably, our model effectively reflects the pattern and style of the dataset, attributed to the capability of encoder-based PLMs. Specifically, as reported in Table 2-ParaDetox, our model excels at capturing semantic information with the highest MAUVE score of 0.733, despite ParaDetox being a challenging colloquial dataset including slang, numerous abbreviations, and various perturbations. In addition, our model demonstrate robustness to such perturbations with 69.5 PPL, while GPT-2 shows a lower performance of 389.1 PPL. Diffusion-EAGS demonstrates high level of text quality surpassing that of GPT-2 in Table 3 in text quality.
As for diversity, our model consistently outperforms GPT models in VS(ngram) and VS(emb) in Table 23 and 4. These results underscore our model’s efficacy in generating diverse responses. Specifically, in the context of story generation as shown in Table 4, our model not only demonstrates effectiveness in enhancing both the quality and diversity of the text but also maintains coherence with the contextual storyline. This is further supported by the high LLM-c scores, which confirm the plausibility of the narratives produced by our model.
D.2 Ours vs CDLMs
Notably, CDLMs demonstrate a noticeable deficiency in diversity. In contrast, our model excels at producing significantly more diverse sequences. This enhanced performance can be attributed to the integration of PLM into our framework, which evidently enriches the diversity of the generated outputs. In addition, our model generally achieves higher text quality scores than CDLMs across all experiments. In case, our model shows a lower PPL than DINOISER, but DINOISER’s MAUVE score approaches zero, indicating that their generated responses are not based on the dataset. Diffusion-EAGS not only surpasses Diffuseq and DINOISER in PPL and SOME but also achieves a MAUVE score improvement of over 0.4, thus indicating that the generated responses accurately reflect the dataset while preserving grammatical integrity. Furthermore, our models require only a few steps, resulting in a 40% reduction of costs with higher quality and diversity.
D.3 Ours vs DDLMs
Contrary to CDLMs, DiffusionBERT exhibits limitations in text quality, as shown in Table 234. On the other hand, our model achieves significantly higher scores across all quality metrics. This observation suggests that the naive application of BERT (Devlin et al., 2019) within the diffusion process fails to fully harness the capabilities of PLMs. SEDD, an open-ended generation model shows low performance under text generation in highly constrained tasks such as ParaDetox. Even in open-ended generation settings of Table 4, Diffusion-EAGS shows higher text quality and semantic scores. Additionally, our model does not require more than 5 diffusion steps, as it can adaptively recover the absorbing state through an entropy-based generation approach. Above results further exemplify the high performance and the time efficiency of our model.
Appendix E Experimental Outputs
E.1 LLM Evaluation
@@ E.2 Tables
Appendix F Well-Generated Output Examples
F.1 ParaDetox